# Efficient Verification-Based Face Identification

### v1

@@ I Introduction
In this work, we suggest an alternative solution in which a hypernetwork [6] $h$ is trained to generate, given a single image $x^{i}_{enroll}$ of a person $i$, the weights $\theta^{i}$ of a verification network $f(\theta^{i})$ that identifies the specific individual ii. Our model is trained using weighted binary cross-entropy loss, and we introduce a new method for effective hard negative mining. The new scheme, termed $K$-means centered sampling (KCS), helps our model learn a more discriminative function by constructing batches containing similar identities. We benchmark our method using six publicly available datasets and demonstrate that it can dramatically reduce the number of parameters while maintaining highly competitive results.
@@ II Background
Deep learning is the main driver behind modern face recognition systems. Many systems are based on a deep neural network $h_{bb}$, which projects an input face image $x_{enroll}$ to an embedding vector $e_{enroll}=h_{bb}(x_{enroll})$, where $e_{enroll}\in R^{d}$ and $d$ is the embedding dimension. Face recognition consists of two phases: enrollment and authentication. In the enrollment stage, a face image of a person $i$ is fed to $h_{bb}$, generating a unique feature vector $e^{i}_{enroll}$ based on their unique facial characteristics. $e^{i}_{enroll}$ is then stored in a database $D\in R^{N\times d}$, where NN is the number of faces stored in the database.
The difference between identification and verification tasks can be formally defined by NN, where $N>1$, and $N=1$ are used in identification and verification, respectively. For authentication, a face image $x_{auth}$ is fed to the feature extractor $e_{auth}=h_{bb}(x_{auth})$, where $e_{auth}$ is the embedding vector of $x_{auth}$. For the verification task, $e_{auth}$ is compared to $e^{i}_{enroll}$ using a predefined distance metric, such as the L2 distance, cosine similarity, or similar. Suppose the distance between the two embedding vectors is lower than a threshold. In that case, the face is authenticated as belonging to the same identity, i.e., $x_{auth}$ is an image of person ii. For identification, the authentication is performed in two phases. The initial search is typically performed using a $K$-Nearest-Neighbours [11] algorithm or similar that identifies the closest embedding $e^{nearest}_{enroll}$ and performs verification between $e^{nearest}_{enroll}$ and $e_{auth}$ using the threshold scheme discussed above.
@@ IV-B Inference Phase
The inference phase describes the actions of the model when a data stream is fed to it. Our inference phase is conducted on the user edge device and requires only the computational burden of $f(\theta^{i})$. A facial image is streamed to the model for verification. Then, it is preprocessed in the same manner as in the enrollment phase. The preprocessed image $x_{stream}\in\mathbb{R}^{h\times w\times 3}$ is then utilized using $y=f(x_{stream},\theta^{i})$. The output, $y$, is a binary decision on whether $x_{stream}$ is an image of user ii.
Figure 1: Illustration of the inference phase of our method. Enrollment is performed once per user $i$, using a single facial image $x^{i}_{enroll}$. The enrollment output is a set of weights $\theta^{i}$ of $f$, designed to fit low-power devices. In the streaming phase, face images $x_{stream}$ are presented to $f$, which then decides whether those belong to person ii. In the stream phase, $h$ is discarded; thus, the main computational burden is mitigated.
@@ IV-C Training Phase
Training (Fig 3) starts by taking an input batch consisting of nBnB samples from $B$ different identities. Without loss of generality, the following descriptions illustrate the case of balanced sampling, i.e., the number of samples per identity $n$ is constant across the batch. This removes bias towards a specific individual within the batch. Our method also applies to unbalanced sampling, where the constant $n$ can be replaced by a per-user $n^{i}$. Let us first define an input training batch $X=\{x^{1}_{1},..,x^{1}_{n},x^{2}_{1},..,x^{2}_{n},..,x^{B}_{1},..x^{B}_{n}\}\in\mathbb{R}^{n\times B\times 3\times h\times w}$, where the superscript represents the identity and the subscript denotes the sample index in the batch running from 1 to $n$. At each training iteration, the goal of $h(X)$ is to generate a weights tensor $\vec{\theta}$, which is $n\times B$ concatenated weight sets corresponding to the input images. Each sample in $j\in X$ results in a weight set $\theta_{j}$ such that $f(x,\theta_{j})$ will recognize only the images of the same identity as the identity of image $j$. Let us define an identity mapper that inputs a sample index $j$ and outputs an identity index, formally: $M:j\in[0,nB)\rightarrow i$. Next, we can define the target matrix $Y\in\mathbb{R}^{nB\times nB}$, where each row represents the target for sample $j$ as follows
@@
TABLE I: Accuracy results on six leading face verification datasets. We present the top-1 accuracy results, the number of parameters of the inference model, and the number of parameters of the model responsible for the user’s enrollment. An identical number of parameters indicates that the enrollment and inference models are identical. We further show the MFLOPS (million floating point operations) for the inference step.
@@ V Experiments
Mini Batch of NNs On a mini-batch of data Some works that train hypernetworks [53] use an inner optimization loop to optimize the parameters $\theta$ for $f$ and use the result as supervision for $h_{gen}$. The downside of this methodology is that, currently, no framework supports backward propagation on a batch of neural networks. This forces a loop over all the batch items and optimization over $\theta_{i}\forall i$ in the batch. As described in this work, the proposed recipe is different in that it relies on the final classification signal. This is because $f$ is a tiny neural network that reduces the concern of vanishing gradients. Nevertheless, at each forward pass, we need to make a forward and a backward pass-through nBnB neural networks $f(\theta)$, with each pass containing nBnB samples. Looping over nBnB neural networks would make the training procedure almost impractical; thus, we would like to tensorize this procedure. This requires using the functorch library [54], which enables forward propagation on a batch of $(nB)^{2}$ images through a batch of nBnB neural networks, resulting in a fast and efficient training regime.
@@ V-A Results
Table LABEL:tab:main_results presents top-1 accuracy results on the six leading face verification datasets mentioned above. Although our method requires a large enrollment model, this is used only once per user, and only the inference model is deployed on the edge device. When comparing inference model size, i.e., its number of parameters and amount of MFLOPs required, we can see the significant advantage of our model. Although our model is almost 22 times smaller than its smallest competitor [24], it maintains on par or better results on all available datasets when the two are compared. Even compared to more than 100 times larger models, we still obtain state-of-the-art results on the CFP-FP dataset. For the first time, we demonstrate a method that dramatically reduces memory requirements while maintaining reasonable results on multiple benchmarks. Our method is based on a state-of-the-art face recognition backbone for the initial embedding during enrollment. It is improved when using the more accurate backbone (AdaFace [17] instead of FaceTransformer[50]). Still, the improvement is slight, and both backbones are comparable. A visual overview of the results on the five datasets: LFW, CA-LFW, CP-LFW, CFP-FP, and AgeDB-30, is shown in Fig. 4, only the methods that have results on all five datasets are presented. Fig. 4 shows both the top-1 accuracy vs. the number of parameters and the number of floating point operations in mega-flops. We expect further improvements when using backbones trained on cleaner and larger datasets, such as subsets of the WebFace260M dataset [33], similarly to the improvement obtained by Kim et al. [17] when training on larger datasets.
@@ VI Discussion
Third, previous works in generative models showed that methods that generate with iterative updates [62](cite diffusion model works) are better than one-step generation models, such as GANs [63]. This suggests that generating the $\theta^{i}$ in a few steps, for example, using a diffusion process, [64] may be beneficial.

### v2

@@ I Introduction
In this work, we suggest an alternative solution in which a hypernetwork [6] $h$ is trained to generate, given a single image $x^{i}_{enroll}$ of a person $i$, the weights $\theta^{i}$ of a verification network $f(\theta^{i})$ that identifies the specific individual $i$. Our model is trained using weighted binary cross-entropy loss, and we introduce a new method for effective hard negative mining. The new scheme, termed $K$-means centered sampling (KCS), helps our model learn a more discriminative function by constructing batches containing similar identities. We benchmark our method using six publicly available datasets and demonstrate that it can dramatically reduce the number of parameters while maintaining highly competitive results.
@@ II Background
Deep learning is the main driver behind modern face recognition systems. Many systems are based on a deep neural network $h_{bb}$, which projects an input face image $x_{enroll}$ to an embedding vector $e_{enroll}=h_{bb}(x_{enroll})$, where $e_{enroll}\in R^{d}$ and $d$ is the embedding dimension. Face recognition consists of two phases: enrollment and authentication. In the enrollment stage, a face image of a person $i$ is fed to $h_{bb}$, generating a unique feature vector $e^{i}_{enroll}$ based on their unique facial characteristics. $e^{i}_{enroll}$ is then stored in a database $D\in R^{N\times d}$, where $N$ is the number of faces stored in the database.
The difference between identification and verification tasks can be formally defined by $N$, where $N>1$, and $N=1$ are used in identification and verification, respectively. For authentication, a face image $x_{auth}$ is fed to the feature extractor $e_{auth}=h_{bb}(x_{auth})$, where $e_{auth}$ is the embedding vector of $x_{auth}$. For the verification task, $e_{auth}$ is compared to $e^{i}_{enroll}$ using a predefined distance metric, such as the L2 distance, cosine similarity, or similar. Suppose the distance between the two embedding vectors is lower than a threshold. In that case, the face is authenticated as belonging to the same identity, i.e., $x_{auth}$ is an image of person $i$. For identification, the authentication is performed in two phases. The initial search is typically performed using a $K$-Nearest-Neighbours [11] algorithm or similar that identifies the closest embedding $e^{nearest}_{enroll}$ and performs verification between $e^{nearest}_{enroll}$ and $e_{auth}$ using the threshold scheme discussed above.
@@ IV-B Inference Phase
The inference phase describes the actions of the model when a data stream is fed to it. Our inference phase is conducted on the user edge device and requires only the computational burden of $f(\theta^{i})$. A facial image is streamed to the model for verification. Then, it is preprocessed in the same manner as in the enrollment phase. The preprocessed image $x_{stream}\in\mathbb{R}^{h\times w\times 3}$ is then utilized using $y=f(x_{stream},\theta^{i})$. The output, $y$, is a binary decision on whether $x_{stream}$ is an image of user $i$.
Figure 1: Illustration of the inference phase of our method. Enrollment is performed once per user $i$, using a single facial image $x^{i}_{enroll}$. The enrollment output is a set of weights $\theta^{i}$ of $f$, designed to fit low-power devices. In the streaming phase, face images $x_{stream}$ are presented to $f$, which then decides whether those belong to person $i$. In the stream phase, $h$ is discarded; thus, the main computational burden is mitigated.
@@ IV-C Training Phase
Training (Fig 3) starts by taking an input batch consisting of $nB$ samples from $B$ different identities. Without loss of generality, the following descriptions illustrate the case of balanced sampling, i.e., the number of samples per identity $n$ is constant across the batch. This removes bias towards a specific individual within the batch. Our method also applies to unbalanced sampling, where the constant $n$ can be replaced by a per-user $n^{i}$. Let us first define an input training batch $X=\{x^{1}_{1},..,x^{1}_{n},x^{2}_{1},..,x^{2}_{n},..,x^{B}_{1},..x^{B}_{n}\}\in\mathbb{R}^{n\times B\times 3\times h\times w}$, where the superscript represents the identity and the subscript denotes the sample index in the batch running from 1 to $n$. At each training iteration, the goal of $h(X)$ is to generate a weights tensor $\vec{\theta}$, which is $n\times B$ concatenated weight sets corresponding to the input images. Each sample in $j\in X$ results in a weight set $\theta_{j}$ such that $f(x,\theta_{j})$ will recognize only the images of the same identity as the identity of image $j$. Let us define an identity mapper that inputs a sample index $j$ and outputs an identity index, formally: $M:j\in[0,nB)\rightarrow i$. Next, we can define the target matrix $Y\in\mathbb{R}^{nB\times nB}$, where each row represents the target for sample $j$ as follows
@@
TABLE I: Accuracy results on six leading face verification datasets. We present the verification accuracy results, the number of parameters of the inference model, and the number of parameters of the model responsible for the user’s enrollment. An identical number of parameters indicates that the enrollment and inference models are identical. We further show the MFLOPS (million floating point operations) for the inference step.
@@ V Experiments
Mini Batch of NNs On a mini-batch of data Some works that train hypernetworks [53] use an inner optimization loop to optimize the parameters $\theta$ for $f$ and use the result as supervision for $h_{gen}$. The downside of this methodology is that, currently, no framework supports backward propagation on a batch of neural networks. This forces a loop over all the batch items and optimization over $\theta_{i}\forall i$ in the batch. As described in this work, the proposed recipe is different in that it relies on the final classification signal. This is because $f$ is a tiny neural network that reduces the concern of vanishing gradients. Nevertheless, at each forward pass, we need to make a forward and a backward pass-through $nB$ neural networks $f(\theta)$, with each pass containing $nB$ samples. Looping over $nB$ neural networks would make the training procedure almost impractical; thus, we would like to tensorize this procedure. This requires using the functorch library [54], which enables forward propagation on a batch of $(nB)^{2}$ images through a batch of $nB$ neural networks, resulting in a fast and efficient training regime.
@@ V-A Results
Table LABEL:tab:main_results presents verification accuracy results on the six leading face verification datasets mentioned above. Although our method requires a large enrollment model, this is used only once per user, and only the inference model is deployed on the edge device. When comparing inference model size, i.e., its number of parameters and amount of MFLOPs required, we can see the significant advantage of our model. Although our model is almost 22 times smaller than its smallest competitor [24], it maintains on par or better results on all available datasets when the two are compared. Even compared to more than 100 times larger models, we still obtain state-of-the-art results on the CFP-FP dataset. For the first time, we demonstrate a method that dramatically reduces memory requirements while maintaining reasonable results on multiple benchmarks. Our method is based on a state-of-the-art face recognition backbone for the initial embedding during enrollment. It is improved when using the more accurate backbone (AdaFace [17] instead of FaceTransformer[50]). Still, the improvement is slight, and both backbones are comparable. A visual overview of the results on the five datasets: LFW, CA-LFW, CP-LFW, CFP-FP, and AgeDB-30, is shown in Fig. 4, only the methods that have results on all five datasets are presented. Fig. 4 shows both the verification accuracy vs. the number of parameters and the number of floating point operations in mega-flops. We expect further improvements when using backbones trained on cleaner and larger datasets, such as subsets of the WebFace260M dataset [33], similarly to the improvement obtained by Kim et al. [17] when training on larger datasets.
@@ VI Discussion
Third, previous works in generative models showed that methods that generate with iterative updates [62] are better than one-step generation models, such as GANs [63]. This suggests that generating the $\theta^{i}$ in a few steps, for example, using a diffusion process, [64] may be beneficial.